# NM-FlowGAN: Modeling sRGB Noise with a Hybrid Approach based on Normalizing Flows and Generative Adversarial Networks

### v2

@@ 3.2 Hybrid Architecture
3.2.1 Pixel-wise Noise Modeling Network
In our pixel-wise noise modeling network, the distribution of each pixel is assumed to be identical, and Gaussian distributions are estimated for each pixel. Our network, based on Normalizing Flows, comprises three flow layers: conditional linear flow, signal-dependent conditional linear flow, and structure-aware conditional linear flow. Each layer effectively estimates the pixel-wise noise distribution by considering the factors affecting it.
Recently, a simple invertible form of the conditional linear flow layer has been proposed in sRGBFlow. This layer linearly transforms the input feature using values determined by the camera type $\delta$ and ISO setting $\gamma$:
@@ 3.2.1 Pixel-wise Noise Modeling Network
where $z^{(l)}$ denotes the input of the $i$th pixel in the $l$th flow layer, $\otimes$ denotes element-wise multiplication, and $F_{cs}(\cdot)$ and $F_{ct}(\cdot)$ are simple neural networks designed to output the scale and translation factors, considering the values of $\delta$ and $\gamma$. While this layer can perform a transformation that considers the camera conditions, the non-linearity and complexity of real sRGB noise makes it difficult to model accurately with such a simple transformation.
As an extension of this work, we designed layers to estimate the pixel-wise noise distribution, considering factors that affect it, such as the clean image intensity and the image structure. First, the signal-dependent conditional linear flow layer performs a linear transformation, considering the clean image intensity:
@@
where $F_{ss}(\cdot)$ and $F_{st}(\cdot)$ consist of pointwise convolution layers and non-linear activation layers, designed to output scale and translation factors based on the clean image intensity and camera conditions. Another flow layer we designed is the structure-aware conditional linear flow layer, which performs a linear transform by considering the surrounding structure of pixel $i$:
@@
where $F_{ts}(\cdot)$ and $F_{tt}(\cdot)$ are designed to have limited receptive fields with $3\times 3$ convolution layers and non-linear activation layers. They output scale and translation factors, considering both the image structure and camera conditions.
@@
3.2.2 Spatial Correlation Modeling Network
Given the synthesized pixel-wise noise, our GAN-based spatial correlation modeling network models the pixel-to-pixel correlation to reduce the discrepancy between the real noise $n$ and synthesized pixel-wise noise $n^{\prime}$. Specifically, as represented in Eq. 5, the final noise value is determined by considering the surrounding pixel-wise noise values. We adopt the U-Net [30] for the generator network and adopt a VGG network-based architecture [31] for the discriminator network. Consequently, by using our spatial correlation modeling network $G(\cdot)$, we can generate the final synthesized sRGB noise image $\tilde{n}$ as follows:
@@ 3.2.2 Spatial Correlation Modeling Network
Fig. 4 provides visual examples of a pixelwise noisy image and a noisy image considering spatial correlation.
Figure 4: Visualization of sRGB noisy image synthesis. (a) clean image $x$, (b) synthesized pixel-wise noisy image $n^{\prime}+x$, and (c) final synthesized noisy image considering spatial correlation $\tilde{n}+x$, (d) real noisy image $y$.
3.2.3 Loss Functions
NLL Loss.
Our pixel-wise noise modeling network $F$ is composed of a sequence of $L$ flow layers:
@@ 3.2.3 Loss Functions
For the sake of clarity, we define $z^{(0)}\stackrel{{\scriptstyle\text{def}}}{{=}}n$ and $z^{(L)}\stackrel{{\scriptstyle\text{def}}}{{=}}z$ for conciseness. To train this network, we employ the NLL loss. It can be formulated as:
$L_{NLL}(n|x,\delta,\gamma)=-p_{z}(F(n|x,\delta,\gamma))-\sum_{l=0}^{L-1}log|det(\mathcal{D}F^{(l)}(z^{(l)}|x,\delta,\gamma))|,$ (13)
@@
GAN Loss.
To train our generator network $G$, we employ WGAN-GP [3, 14] approach to calculate the adversarial loss. The adversarial loss for $G$ is given by:
@@
where ||||| | represents channel-wise concatenation and $sg(\cdot)$ denotes stop gradient operation [8]. The discriminator, denoted as $D(\cdot)$, evaluates the authenticity of the synthesized noise with consideration of the input clean image. In addition, $\lambda$ is the parameter for the balance between the pixel-wise noise modeling networks and the spatial correlation modeling networks.
4 Experiments
4.1 Experimental Setup
4.1.1 Dataset.
@@ 4.1 Experimental Setup
To evaluate the performance of sRGB noise modeling, we use the Smartphone Image Denoising Dataset (SIDD [2]). The SIDD medium split contains 320 noisy-clean image pairs captured with various camera types: LG G4 (G4), Google Pixel (GP), iPhone 7 (IP), Motorola Nexus 6 (N6), and Samsung Galaxy S6 Edge (S6), each under various ISO settings. For a fair comparison, we follow the sRGBFlow [20] dataset settings, using 80% of the data for training and 20% of the data for validation.
4.1.2 Implementation Details.
@@
We extract patches of size $96\times 96$ with a step size of $48$ to optimize NM-FlowGAN. During training, we augment all of the training patches by randomly flipping and rotating them by $90^{\circ}$. We employ the Adam [18] optimizer for training with an initial learning rate set to $10^{-4}$. The learning rate decreases by half every 10 epochs, and our model trains over 40 epochs. In addition, we set $\lambda$ to $0.5$ in our experiments.
Table 1: KL divergence between synthesized noise and real noise calculated under various camera types. Bold numbers indicate the best performance (lower is better), while underlined numbers indicate the second best. $*{}^{*}$ denotes the method of synthesizing sRGB noise with a paired noisy image, which is a more constrained approach that should be considered when interpreting this table.
Method G4 GP IP N6 S6 ALL
AWGN 0.067 0.114 0.056 0.147 0.084 0.086
Heteroscedastic 0.072 0.139 0.058 0.133 0.082 0.078
C2N [16] 0.230 0.446 0.117 0.240 0.682 0.388
sRGBFlow [20] 0.042 0.067 0.039 0.053 0.044 0.048
NeCA-W [12] 0.089 0.087 0.045 0.085 0.072 0.071
$*{}^{*}$ [12] 0.061 0.026 0.028 0.043 0.034 0.034
Ours 0.039 0.026 0.017 0.041 0.033 0.028
@@
Table 2: Ablation study on different architectural choices. SAL the structure-aware conditional linear flow layer, SDL the signal-dependent conditional linear flow layer, and GAN the spatial correlation network based on GAN.
@@
0.037
0.555
0.128
0.037
0.028
4.1.3 Baselines.
Our NM-FlowGAN is compared with several existing sRGB noise modeling methods in the experiments. We include traditional methods such as Additive White Gaussian Noise(AWGN), heteroscedastic Gaussian noise, and deep learning-based modeling methods such as C2N [16], sRGBFlow [20], NeCA-W [12]. For the synthesis using AWGN and heteroscedastic Gaussian noise, we estimate the parameters corresponding to each noise model and apply them to each pixel. For C2N, sRGBFlow, and NeCA-W, we use the official codes provided by the authors. Specifically, we report two types of performance for NeCA-W.
In the NeCA [12], it is mentioned that NeCA-W uses arbitrary noisy images to estimate the gain map of the target image. However, in their experiments, they use paired noisy images corresponding to the clean images instead of arbitrary noisy images to eliminate randomness in the evaluation. After synthesizing the noise, as a noise modeling performance, the KL divergence of the synthesized noise image from the real noise image extracted from the paired noisy image is calculated. This approach may lead to an overestimation of the noise modeling performance as it considers the ground truth data during the noise synthesis. Moreover, paired noisy images may not be achievable in real-world noise synthesis scenarios. Therefore, in our experiments, we include both experimental results using arbitrary noisy images captured under the same camera conditions of the target clean images, as described in the method section of the NeCA paper (NeCA-W), and experimental results using paired noisy images ($*{}^{*}$) as baselines. In addition, to eliminate the randomness of the evaluation, we evaluate the performance of NeCA-W over 10 iterations and report the average results.
4.2 Results on sRGB Noise Modeling
Tab. 1 lists the discrete Kullback-Leibler(KL) divergence results for NM-FlowGAN and several baseline models. We evaluate the performance of the noise modeling using the KL divergence, which determines the similarity between the distributions of the real and synthesized noise by comparing the histogram. Specifically, the histogram range is set from -260 to 260, with 130 intervals. In our experimental results, NM-FlowGAN outperforms other baseline models. Notably, our NM-FlowGAN shows a KL divergence that is 0.006 lower than $*{}^{*}$ which uses paired noisy images. That is, our method shows a lower KL divergence even without using paired noisy images. Additionally, our method shows a reduction in KL divergence of 0.043 and 0.020 compared to NeCA-W and sRGBFlow. We also provide a qualitative comparison between the real noise images and the synthesized noise generated by our NM-FlowGAN and other baselines in Fig. 5.
4.3 Ablation Study
Tab. 2 summarizes the KL divergence for different architectural choices of our proposed NM-FlowGAN. In the experiment using only a GAN-based architecture, we observe a high KL divergence due to the lack of consideration for features that affect the noise distribution, In the experiments using the flow layers, the KL divergences are gradually decreased with the addition of signal-dependent and structure-aware conditional linear flow layers. Notably, in the experiment consisting only of flow layers (without GAN), it is observed that the KL divergence is marginally higher than when both flow layers and GAN are used simultaneously. This is because KL divergence, which is based on histogram similarity, may not effectively capture factors that are not revealed in the histogram, such as spatial correlation. Therefore, in the following subsection, we further evaluate the closeness of the synthesized noise to the real noise by comparing the performance of image denoising networks trained with the synthesized noise from each method.
Table 3: Image denoising results on the SIDD benchmark dataset. Each denoiser is trained on synthesized noisy-clean image pairs generated by each model. $*{}^{*}$ denotes the method of synthesizing sRGB noise with a paired noisy image, which is a more constrained approach that should be considered when interpreting this table.
@@ 4.3 Ablation Study
Heteroscedastic 32.13 0.843
C2N [16] 33.76 0.901
@@
Real Noisy 37.73 0.941
Table 4: SIDD validation results under different architectural choices. The meanings of SAL, SDL, and GAN are as described in Tab. 2.
@@
4.4 Results on Real-world Image Denoising
4.4.1 Preparation.
To further investigate the quality of the synthesized noise, we evaluate its performance in real-world image denoising. We use the widely adopted image denoiser DnCNN [33] to train on the synthesized noisy-clean image pairs generated by each of the baselines. The clean images used to generate the synthesized noisy-clean image pairs are obtained from the sRGB images in the SIDD medium split. In addition, we adopt the SIDD validation and benchmark dataset for validation and testing. The SIDD validation and benchmark dataset contain 1,280 patches, each of size $256\times 256$. Although the ground truth for the SIDD benchmark dataset used in testing is not publicly available, peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) results for denoising can be obtained through the online submission system provided on the SIDD website [^1] https://www.eecs.yorku.ca/~kamel/sidd/benchmark.php. In our real-world image denoising experiments, we include not only the baselines used in sRGB noise modeling but also a supervised method trained on real noisy-clean image pairs to estimate the upper bound performance.
4.4.2 Results.
@@ 4.4 Results on Real-world Image Denoising
Tab. 3 lists the performance of an image denoising networks trained on noisy-clean image pairs synthesized by our proposed method in comparison to other baselines. The results indicate that our method outperforms the other baselines in terms of PSNR and SSIM. Specifically, NM-FlowGAN not only achieves significant PSNR gains of 1.11dB and 2.30dB over NeCA-W and sRGBFlow respectively, but also outperforms the more constrained $*{}^{*}$ by 0.22dB. These results indicate that the noise generated by our model has a more similar distribution to the real noise compared to the other baselines. A visual comparison between NM-FlowGAN and the baselines is shown in Fig. 6.
Figure 6: Visual comparison of denoised images in the SIDD validation dataset. Each denoiser is trained on synthesized noisy-clean image pairs. Best view in zoom-in.
4.4.3 Ablation Study.
@@
Tab. 4 summarizes the image denoising performance of different architecture choices for our method. Notably, the experiment consisting only of flow layers (without GAN), while showing a low KL divergence in Tab. 2, shows a dramatically degraded image denoising performance. This indicates that efficient modeling of spatial correlation plays a crucial role in improving the applicability of the noise model. In experiments using the flow layers based on the hybrid architecture (with GAN), both PSNR and SSIM are progressively improved with the addition of signal-dependent and structure-aware conditional linear flow layers.
5 Conclusion
In this paper, we introduce a novel hybrid method for modeling sRGB noise. We begin by analyzing the characteristics of sRGB noise and exploring the factors that affect its distribution. From this analysis, we present NM-FlowGAN, a noise modeling neural networks that combines two generative neural networks: Normalizing Flows and GAN. By exploiting the strengths of both networks, our method demonstrates a robust performance to effectively model sRGB noise. Our experimental results show that our method outperforms other baseline methods in sRGB noise modeling. Furthermore, we show that image denoising neural networks, when trained on noisy-clean image pairs synthesized by our model, outperform those trained with other existing noise models.
6 Losses for Adversarial Networks
In this section, we describe the details of training spatial correlation modeling networks based on generative adversarial networks (GAN). To facilitate the training of the network in an adversarial manner, we adopt Wasserstein GAN with gradient penalty (WGAN-GP) algorithm [14]. This algorithm is adopted to improve training stability and prevent model collapse. WGAN-GP achieves this by not only defining adversarial loss functions for the generator and the discriminator (called the critic in WGAN), but also introducing a gradient penalty term specifically for the discriminator.
As mentioned in the main paper, the adversarial loss for our generator $G$ is expressed as:
$L_{adv}(x,n^{\prime})=-\lambda\cdot D(x||G(sg(n^{\prime}))),$ (15)
where $D(\cdot)$ denotes our discriminator, and $x$ and $n^{\prime}$ denote the clean image and the synthesized pixel-wise noise image, respectively. In addition, $\lambda$ is the balance parameter between the pixel-wise noise modeling networks and the spatial correlation modeling networks, the operator ||||| | denotes the channel-wise concatenation, and $sg(\cdot)$ denotes the stop-gradient operation [8]. This equation indicates that our discriminator evaluates the authenticity of the noise synthesized by $G$, with consideration of the clean image $x$.
To train the discriminator, the critic loss $L_{critic}$ for our discriminator is defined as follows:
$L_{critic}=\lambda\cdot(L_{wgan}+\alpha L_{gp}),$ (16)
$L_{wgan}(x,n,n^{\prime})=D(x||G(sg(n^{\prime})))-D(x||n).$ (17)
In addition, $L_{gp}$ is the gradient penalty loss, defined as :
$L_{gp}(x,n,n^{\prime})=(||\nabla_{\hat{x}}D(\hat{x})||_{2}-1)^{2},$ (18)
7 Simultaneous Training

### v3

@@ III-B Hybrid Architecture
III-C Pixel-wise Noise Modeling Network
In our pixel-wise noise modeling network, the distribution of each pixel is assumed to be identical, and Gaussian distributions are estimated for each pixel. Our network, based on Normalizing Flows, comprises three flow layers: conditional linear flow, signal-dependent conditional linear flow (SDL), and structure-aware conditional linear flow (SAL). Each layer effectively estimates the pixel-wise noise distribution by considering the factors affecting it.
Recently, a simple invertible form of the conditional linear flow layer has been proposed in sRGBFlow [13]. This layer linearly transforms the input feature using values determined by the camera type $\delta$ and ISO setting $\gamma$:
@@ III-C Pixel-wise Noise Modeling Network
where $z^{(l)}$ denotes the input of the $i$th pixel in the $l$th flow layer, $\otimes$ denotes element-wise multiplication, and $F_{cs}(\cdot)$ and $F_{ct}(\cdot)$ are simple neural networks designed to output the scale and translation factors, considering the values of $\delta$ and $\gamma$. While this layer can perform a transformation that considers the camera conditions, the non-linearity and complexity of real sRGB noise make it difficult to model accurately with such a simple transformation.
To address these limitations, we designed enhanced layers that estimate the pixel-wise noise distribution by considering factors such as the clean image intensity and the image structure. SDL performs a linear transformation based on the clean image intensity:
@@
where $F_{ss}(\cdot)$ and $F_{st}(\cdot)$ consist of pointwise convolution layers and non-linear activation layers, designed to output scale and translation factors based on the clean image intensity and camera conditions. SAL performs a linear transformation by considering the surrounding structure of pixel $i$:
@@
where $F_{ts}(\cdot)$ and $F_{tt}(\cdot)$ are designed to have limited receptive fields with $3\times 3$ convolution layers and non-linear activation layers. They output scale and translation factors considering both the image structure and camera conditions.
Figure 5 illustrates the detailed architecture of the core components of our pixel-wise noise modeling network: SDL and SAL. These layers are designed to estimate and apply scale and bias values for each pixel using a clean image and conditions, including camera type and ISO setting. SDL and SAL first apply a one-hot encoding of the camera type and the ISO setting, and then extract the features using an architecture consisting of residual blocks. This method of encoding camera conditions is adopted from sRGBFlow [13]. The extracted features are then channel-wise concatenated with the clean image to serve as input to the feature extraction networks (FEN) of SDL and SAL. Each of these layers has a different architecture for feature extraction, and the outputs from these layers are used as scale and bias parameters for each pixel.
Specifically, SDL consists mainly of pointwise convolution layers, instance normalization layers, and activation layers. This design aims to estimate log scale and bias parameters based on the clean image intensity of each pixel. On the other hand, SAL consists mainly of convolution layers with $3\times 3$ filters and activation layers. Its design aims to estimate log scale and bias parameters based on the image structure surrounding each pixel.
@@
Figure 6: Visualization of sRGB noisy image synthesis. (a) clean image $x$, (b) real noisy image $y$, (c) synthesized pixel-wise noisy image $n^{\prime}+x$, and (d) final synthesized noisy image considering spatial correlation $\tilde{n}+x$.
III-D Spatial Correlation Modeling Network
Given the synthesized pixel-wise noise, our GAN-based spatial correlation modeling network models the pixel-to-pixel correlation to reduce the discrepancy between the real noise $n$ and synthesized pixel-wise noise $n^{\prime}$. Specifically, as represented in (8), the final noise value is determined by considering the surrounding pixel-wise noise values. We adopt the U-Net [30] for the generator network and a VGG network-based architecture [31] for the discriminator network. Consequently, by using our spatial correlation modeling network $G(\cdot)$, we can generate the final synthesized sRGB noise image $\tilde{n}$ as follows:
@@ III-D Spatial Correlation Modeling Network
Figure 6 provides visual examples of a pixel-wise noisy image and a noisy image considering spatial correlation.
TABLE I: KL divergence between synthesized noise and real noise calculated under various camera types. $\dagger$ denotes the method of synthesizing sRGB noise with a paired noisy image. In addition, G4, GP, IP, S6, and N6 represent the average KL divergence for the respective camera types, and AVG represents the average KL divergence of the entire test set.
Method G4 GP IP N6 S6 AVG
Traditional Methods AWGN 0.0672 0.1136 0.0559 0.1467 0.0836 0.0856
Heteroscedastic 0.0718 0.1393 0.0575 0.1330 0.0823 0.0779
Learning-based Methods (with paired image) NeCA-W [14] 0.0611 0.0260 0.0284 0.0435 0.0337 0.0344
NAFlow [21] 0.0186 0.0681 0.0129 0.0423 0.0164 0.0286
Learning-based Methods (without paired image) C2N [15] 0.2296 0.4460 0.1172 0.2399 0.6822 0.3878
NeCA-W [14] 0.0893 0.0867 0.0452 0.0847 0.0716 0.0706
sRGBFlow [13] 0.0424 0.0669 0.0390 0.0528 0.0439 0.0477
Ours 0.0389 0.0259 0.0164 0.0407 0.0334 0.0283
III-E Loss Functions
NLL Loss: Our pixel-wise noise modeling network $F$ is composed of a sequence of $L$ flow layers:
@@ III-E Loss Functions
without the necessity of training separate neural networks for each camera or using paired noisy images. For the sake of clarity, we define $z^{(0)}\stackrel{{\scriptstyle\text{def}}}{{=}}n$ and $z^{(L)}\stackrel{{\scriptstyle\text{def}}}{{=}}z$ for conciseness. To train this network, we employ the NLL loss. It can be formulated as:
$L_{NLL}(n|x,\delta,\gamma)=-p_{z}(F(n|x,\delta,\gamma))\\ -\sum_{l=0}^{L-1}log|det(\mathcal{D}F^{(l)}(z^{(l)}|x,\delta,\gamma))|,$ (16)
@@
GAN Loss: To train our generator network $G$, we employ WGAN-GP [32, 33] approach to calculate the adversarial loss. The adversarial loss for $G$ is given by:
@@
where $D(\cdot)$ denotes our discriminator, and $x$ and $n^{\prime}$ denote the clean image and the synthesized pixel-wise noise image, respectively. In addition, $\lambda$ is the balance parameter between the pixel-wise noise modeling networks and the spatial correlation modeling networks, the operator ||||| | denotes the channel-wise concatenation, and $sg(\cdot)$ denotes the stop-gradient operation [34]. This equation indicates that our discriminator evaluates the authenticity of the noise synthesized by $G$, with consideration of the clean image $x$. Meanwhile, to ensure that the discriminator is effectively trained with the generator, we define the critic loss $L_{critic}$ for our discriminator as follows:
$L_{critic}=\lambda\cdot(L_{wgan}+\alpha L_{gp}),$ (18)
$L_{wgan}(x,n,n^{\prime})=D(x||G(sg(n^{\prime})))-D(x||n).$ (19)
In addition, $L_{gp}$ is the gradient penalty loss, defined as:
$L_{gp}(x,n,n^{\prime})=(||\nabla_{\hat{x}}D(\hat{x})||_{2}-1)^{2},$ (20)
IV Experiments
IV-A Experimental Setup
@@
Dataset: To evaluate the performance of sRGB noise modeling, we use the Smartphone Image Denoising Dataset (SIDD [1]). The SIDD medium split contains 320 noisy-clean image pairs captured with various camera types: LG G4 (G4), Google Pixel (GP), iPhone 7 (IP), Motorola Nexus 6 (N6), and Samsung Galaxy S6 Edge (S6), each under various ISO settings. For a fair comparison, we follow the sRGBFlow [13] dataset settings, using 80% of the data for training and 20% of the data for validation.
@@
Implementation Details: We extract patches of size $96\times 96$ with a step size of $48$ to optimize NM-FlowGAN [^2] Our code is available at: https://github.com/YoungJooHan/NM-FlowGAN. During training, we augment all of the training patches by randomly flipping and rotating them by $90^{\circ}$. We employ the Adam [35] optimizer for training with an initial learning rate set to $10^{-4}$. The learning rate decreases by half every 10 epochs, and our model trains over 40 epochs. In addition, we set $\lambda$ to $0.5$ in our experiments.
Baselines: Our NM-FlowGAN is compared with several existing sRGB noise modeling methods in the experiments. We include traditional methods such as Additive White Gaussian Noise(AWGN), heteroscedastic Gaussian noise, and deep learning-based modeling methods such as C2N [15], sRGBFlow [13], NeCA-W [14], and NAFlow [21]. For the synthesis using AWGN and heteroscedastic Gaussian noise, we estimate the parameters corresponding to each noise model and apply them to each pixel. For C2N, sRGBFlow, NeCA-W, and NAFlow, we use the official codes provided by the authors. Specifically, $\dagger$ in Table I and Figure 7 refers to a method that uses paired noisy images for noise synthesis.
Using paired noisy images allows for more accurate noise generation. However, the need for paired noisy images for noise synthesis does not take one of the main advantages of noise modeling mentioned in Section 1, which is the ability to generate noisy-clean image pairs in scenes that are not achievable in real-world environments. Therefore, its applicability in real-world scenarios is limited. For NeCA-W, since the method can perform noise synthesis using both paired noisy images and unpaired noisy images, we present the experimental results for both scenarios.
IV-B Results on sRGB Noise Modeling
Table I lists the discrete Kullback-Leibler(KL) divergence results for NM-FlowGAN and several baseline models. We evaluate the performance of the noise modeling using the KL divergence, which determines the similarity between the distributions of the real and synthesized noise by comparing the histogram. Specifically, the histogram range is set from -260 to 260, with 130 intervals.
In our experimental results, our NM-FlowGAN showed the best performance among the methods that do not use paired noisy images for noise synthesis. Specifically, NM-FlowGAN shows a lower KL divergence of 0.043 and 0.020 compared to the recently proposed NeCA-W and sRGBFlow, respectively. Notably, our NM-FlowGAN achieves a slightly lower KL divergence compared to NeCA-W and NAFlow, both of which use paired noisy images. Figure 7 shows a qualitative comparison between the real noise images and the synthesized noise images generated by our NM-FlowGAN and other baselines.
To further validate the noise modeling performance of our model, we present additional experimental results in this section. First, Figure 8 shows the standard deviation as a function of clean image intensity for our NM-FlowGAN, sRGBFlow, and NeCA-W. This figure indicates that our model shows similar behavior to real noise compared to the other baselines.
IV-C Ablation Study
Table II summarizes the KL divergence for different architectural choices of our proposed NM-FlowGAN. In the experiment using only a GAN-based architecture, we observe a high KL divergence due to the lack of consideration for features that affect the noise distribution. In the experiments using the flow layers, the KL divergences are gradually decreased with the addition of signal-dependent and structure-aware conditional linear flow layers. Notably, in the experiment consisting only of flow layers (without GAN), it is observed that the KL divergence is marginally higher than when both flow layers and GAN are used simultaneously. This is because KL divergence, which is based on histogram similarity, may not effectively capture factors that are not revealed in the histogram, such as spatial correlation. Therefore, in the following subsection, we further evaluate the closeness of the synthesized noise to the real noise by comparing the performance of image denoising networks trained with the synthesized noise from each method.
@@ IV-C Ablation Study
TABLE II: Ablation study on different architectural choices. SAL the structure-aware conditional linear flow layer, SDL the signal-dependent conditional linear flow layer, and GAN the spatial correlation network based on GAN.
@@
0.0377
0.5546
0.1276
0.0371
0.0283
TABLE III: Image denoising results on the SIDD benchmark dataset. Each denoiser is trained on synthesized noisy-clean image pairs generated by each model. denotes the method of synthesizing sRGB noise with a paired noisy image.
@@
Learning-based (with paired image) NeCA-W [14] 36.82 0.932
NAFlow [21] 37.22 0.935
Learning-based (without paired image) C2N [15] 33.76 0.901
@@
IV-D Results on Real-world Image Denoising
Preparation: To further investigate the quality of the synthesized noise, we evaluate its performance in real-world image denoising. We use the widely adopted image denoiser DnCNN [36] to train on the synthesized noisy-clean image pairs generated by each of the baselines. The clean images used to generate the synthesized noisy-clean image pairs are obtained from the sRGB images in the SIDD medium split. In addition, we adopt the SIDD validation and benchmark dataset for validation and testing of image denoising networks. The SIDD validation and benchmark dataset contain 1,280 patches, each of size $256\times 256$. Although the ground truth for the SIDD benchmark dataset used in testing is not publicly available, peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) results for denoising can be obtained through the online submission system provided on the SIDD website [^3] https://www.eecs.yorku.ca/~kamel/sidd/benchmark.php. In our real-world image denoising experiments, we include not only the baselines used in sRGB noise modeling but also a supervised method trained on real noisy-clean image pairs to estimate the upper bound performance.
Figure 9: Visual comparison of denoised images in the SIDD validation dataset. Each denoiser is trained on synthesized noisy-clean image pairs generated by each model. Best view in zoom-in.
TABLE IV: SIDD validation results under different architectural choices. The meanings of SAL, SDL, and GAN are as described in Table II.
@@ IV-D Results on Real-world Image Denoising
Results: Table III lists the performance of image denoising networks trained on noisy-clean image pairs synthesized by our proposed method in comparison to other baselines. The results indicate that our method outperforms the other baselines in terms of PSNR and SSIM. Specifically, NM-FlowGAN achieves significant PSNR gains of 1.11dB and 2.30dB over NeCA-W and sRGBFlow respectively. These results indicate that the noise generated by our model has a more similar distribution to the real noise compared to the other baselines. A visual comparison between NM-FlowGAN and the baselines is shown in Figure 9.
@@
Ablation Study: Table IV summarizes the image denoising performance of different architecture choices for our method. Notably, the experiment consisting only of flow layers (without GAN), while showing a low KL divergence in Table II, shows a degraded image denoising performance. This indicates that efficient modeling of spatial correlation plays a crucial role in improving the applicability of the noise model. In experiments using the flow layers based on the hybrid architecture (with GAN), both PSNR and SSIM are progressively improved with the addition of signal-dependent and structure-aware conditional linear flow layers.
TABLE V: The effect of enlarging the SIDD dataset with synthesized samples from the SIDD+ dataset [37]. $+$ denotes that the training dataset has been enlarged by synthesized noisy-clean image pairs generated using the sRGB noise modeling method, and ”Supervised” refers to training on the SIDD dataset only.
Method PSNR(dB) SSIM
Supervised 35.21 0.902
+ sRGBFlow [13] 34.69 0.893
+ NeCA-W [14] 35.57 0.906
+ Ours 35.75 0.908
IV-E Applicability in Real-world Scenario
As mentioned in Section 1, modeling sRGB noise enables the generation of noisy-clean image pairs across various environments, even in scenarios where obtaining such pairs is challenging due to physical constraints. This allows for training image denoising networks with synthesized pairs that are similar to real usage scenarios, thereby enhancing the generalization ability of the model and ensuring robust performance.
Figure 10: Visual comparison of denoising results on the SIDD+ dataset. Each denoiser is trained on the SIDD dataset enlarged by synthesized noisy-clean image pairs generated using the sRGB noise modeling method, and ”Supervised” refers to training on the SIDD dataset only.
To demonstrate the importance of training with scenes similar to real usage environments, we utilize the SIDD+ [37] dataset, which contains scenes different from those in the SIDD dataset. Specifically, we use our NM-FlowGAN to generate synthetic noisy-clean image pairs from 50% of the SIDD+ dataset, and then add these to the SIDD dataset to train the image denoising networks. Subsequently, we evaluate the performance of the denoiser using the remaining 50% of the SIDD+ dataset. In the experiment, the camera settings for noise synthesizing in the SIDD+ dataset are randomly selected for each image.
Table V and Figure 10 demonstrate that image denoising performance improves when networks are trained on datasets that are enlarged with synthesized noisy-clean image pairs derived from scenes similar to actual usage environments. Specifically, when compared to the results using only the SIDD dataset, enlarging the dataset with synthesized noisy-clean image pairs generated by our NM-FlowGAN results in a PSNR gain of 0.54. Furthermore, our proposed method achieves higher gains in PSNR and SSIM compared to recently proposed methods such as sRGBFlow and NeCA-W. These experimental results highlight the accuracy of our sRGB noise modeling method and emphasize its applicability in real-world scenarios.
IV-F Camera-Specific Noise Modeling
We demonstrated that our NM-FlowGAN effectively models the noise in the SIDD dataset, which consists of images captured by various cameras, by training a single unified model. However, as mentioned, sRGB noise is significantly affected by hardware specifications such as image sensors. In scenarios where the noise modeling data is captured by cameras with widely varying hardware specifications, employing camera-specific training becomes a suitable solution to improve performance.
Table VI demonstrates that our proposed NM-FlowGAN shows promising performance even in camera-specific training. This is due to our hybrid architecture, which exploits the advantages of both Normalizing Flows and GAN, allowing our NM-FlowGAN to be stably trained even with small-sized datasets. Specifically, in our pixel-wise noise modeling networks based on Normalizing Flows, stable training can inherently be achieved even on small-sized datasets. Moreover, our GAN-based spatial correlation modeling networks model spatial correlations from synthesized pixel-wise noise, enabling more consistent modeling compared to other GAN-based noise modeling networks. This consistency leads to more stable training.
TABLE VI: KL divergence of our NM-FlowGAN in camera-specific training. indicates results obtained from camera-specific training. In camera-specific training, each noise modeling networks is trained using only the data that corresponds to each camera type.
Method G4 GP IP N6 S6 ALL
NM-FlowGAN 0.0389 0.0259 0.0164 0.0407 0.0334 0.0283
NM-FlowGAN 0.0306 0.0234 0.0103 0.0581 0.0496 0.0318
IV-G Simultaneous Training
Our NM-FlowGAN consists of two distinct neural networks: pixel-wise noise modeling networks and spatial correlation networks. These networks are based on Normalizing Flows and GAN, respectively. Our method is capable of training multiple networks simultaneously. However, each network is optimized independently according to its respective loss function, ensuring that the optimization of one network does not affect the other networks. In this paper, we indicate this training strategy to simultaneous training. This implementation is possible because each neural network tends to synthesize a distinct form of noise, and it can improve the training stability and performance of our model in sRGB noise synthesis. Table VII shows the comparison of noise modeling performance for three different training strategies: the simultaneous training strategy, the joint training strategy without using a stop gradient function, and the two-stage training strategy where the pixel-wise and spatial correlation modeling networks are trained sequentially.